OpenStudy (anonymous):

for the equation given below, y/x+8y=x^9-4 evaluate dy/dx at point (1,-3/25)

OpenStudy (freckles):

Troubles differentiating I suppose?

OpenStudy (freckles):

\[\frac{y}{x}+8y=x^9-4 \text{ is correct interpretation ?} \]

OpenStudy (anonymous):

sorry i meant (y)/(x+8y)=x^9-4

OpenStudy (anonymous):

8y under the y as well

OpenStudy (freckles):

\[\frac{y}{x+8y}=x^9-4\] I bet you have no problem with the right hand side

OpenStudy (freckles):

The left hand side will require us to know quotient rule

OpenStudy (freckles):

\[\frac{d}{dx}(\frac{y}{x+8y})=\frac{(y)'(x+8y)-y(x+8y)'}{(x+8y)^2}\]

OpenStudy (freckles):

This is only leaves you with differentiating y and (x+8y)

OpenStudy (freckles):

\[\frac{d}{dx}y=\frac{dy}{dx} \text{ (note: which people also call this) } =y'\]

OpenStudy (freckles):

Try finding \[\frac{d}{dx}(x+8y)=?\]

OpenStudy (anonymous):

ok i got, [1+8 d/dx (y(x))\]

OpenStudy (freckles):

ok so we have 1+8y' coolness

OpenStudy (freckles):

ok so let's put all of that in...

OpenStudy (freckles):

\[\frac{y'(x+8y)-y(1+8y')}{(x+8y)^2}=9x^8\]

OpenStudy (freckles):

Now we need to solve for y'...

OpenStudy (freckles):

first I would multiply both side by that ickyness in the bottom on the left side

OpenStudy (freckles):

Then after you have done I would start using distributive property where needed on left. Then gather my terms that y' and put everything else on the other side

OpenStudy (anonymous):

i dont get what you mean by multiply both sides

OpenStudy (anonymous):

multiply both sides by what?

OpenStudy (freckles):

by the ickyness on the bottom on the left

OpenStudy (freckles):

There is a fraction on the left a fraction has a top and a bottom

OpenStudy (freckles):

\[\cancel{(x+8y)^2} \cdot \frac{y'(x+8y)-y(1+8y')}{\cancel{(x+8y)^2}}=9x^8(x+8y)^2\]

OpenStudy (freckles):

I multiplied both sides by the ickyness that was on the bottom of the left

OpenStudy (freckles):

I'm doing this because I'm trying to isolate y'

OpenStudy (freckles):

That is the side with y'

OpenStudy (freckles):

To get to y' I have to undo that division first

OpenStudy (anonymous):

oh ok..

OpenStudy (anonymous):

got you

OpenStudy (anonymous):

now do i distribute on the left?

OpenStudy (freckles):

Yep If you can actually skip that step if you are really good with algebra and sorta like basically do it in your head but yeah

OpenStudy (freckles):

You want to gather your terms that have y' together

OpenStudy (anonymous):

xy'+8yy'-7+8yy'?

OpenStudy (anonymous):

is that correct for the left side?

OpenStudy (freckles):

\[xy'+8yy'-y-8yy'\]

OpenStudy (anonymous):

yah i meant y instead of 7 lol

OpenStudy (freckles):

and a - in front of the 8yy'

OpenStudy (freckles):

but anyways you should see something cancel there

OpenStudy (freckles):

before we move on

OpenStudy (anonymous):

so xy'-y

OpenStudy (freckles):

\[xy'-y=9x^8(x+8y)^2\]

OpenStudy (freckles):

Do you think you can finish solving for y' now?

OpenStudy (freckles):

You only need 2 more steps

OpenStudy (anonymous):

add y and divide by x correct?

OpenStudy (freckles):

yep

OpenStudy (freckles):

\[y'=\frac{9x^8(x+8y)^2+y}{x}\]

OpenStudy (anonymous):

yep now i will plug in the x and y values

OpenStudy (anonymous):

and see if its right

OpenStudy (freckles):

yeah

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

-1.056

OpenStudy (anonymous):

and it is correct

OpenStudy (freckles):

Well I will have to check again I got -.1056 maybe i made mistake then

OpenStudy (anonymous):

thank you freckles! can you help me with another question

OpenStudy (anonymous):

no it is right, i submitted it and it came up correct

OpenStudy (freckles):

oh I thought you put in -1.056

OpenStudy (anonymous):

oh no .-1056

OpenStudy (freckles):

ok cool

OpenStudy (anonymous):

suppose that (g(x))^2+6x=x^2g(x)+17 and g(2) =5 find g'(2)